Find the nth Maclaurin polynomial for the function. - 3x – 3 X + 1 n = 4 PA(X) =
Answer #1

v f(x) = 34 - 3 x 41 110) = of f (x) = 3 (2+1) - (32-3) so - 6 Jex) = 12 po: -12 100= contents 110:3 $08) - Toronto = 1)